If z=sin(x/y) , x=3t , y=5−t^2 dz/dt using the chain rule. Assume the variables are restricted to domains on which the functions are defined.

dz/dt=
Answer #1

given that 7 = sin(ay) k=3t Y= 5-ť de = lesz + on the —4) Waime, and .: 7= sin(nry I Jitterentiate both op del With respect t=3t Jitterentiate both side with respect to t weget = (3+) = 3 y=5- ť Ditterentiate both sides with respect to t Weget Yr